Question 1.

A. With the help of equation show that the aqueous solution of:

i. Ammonia is basic

ii. Hydrogen carbonate is acidic


B. Calculate the pH of the following;

i. 0.01molL⁻1of hydrochloric acid (HCl)

ii. 0.02molL⁻1 of methanoic (HCOOH) acid given Ka (HCOOH) = 1.78 x 10⁻4


C. A pH value can be calculated from [OH⁻] and Kw

i. Calculate the concentration of OH⁻ (aq) in a solution of HCl with a pH of 3.7, given Kw = 1 x 10⁻14

ii. Give formula for the conjugate base of HCl
1
Expert's answer
2020-05-21T13:52:19-0400

Solution.

A.

i. "NH_3+H_2O\to NH_4OH;"

ii. "H_2CO_3+2H_2O\to2H_3O+CO_3;"

B.

i. "[HCl]=0.01molL^{-1};"

"pH=-log[H_3O^+];"

"pH=-log0.01=2;"

ii."[HCOOH]=0.02molL^{-1};"

"K_\alpha(HCOOH)=1.78\sdot10^{-4};"

"K_\alpha=\dfrac{[H_3O^+][HCO_2^-]}{[HCO_2H]};"

"[H_3O^+]=[HCO_2^-]=x;"

"x^2=K_\alpha[HCO_2H];"

"x^2=1.78\sdot10^{-4}\sdot0.02=0.0356\sdot10^{-4}; x=0.189\sdot10^{-2}" ;

"[H_3O^+]=0.189\sdot10^{-2};"

"pH=-log(0.189\sdot10^{-2})=log529=2.72;"

C.

i. "K_b=1\sdot10^{-4};"

"pH=3.7;"

"pH=-log[H_3O];\implies[H_3O^+]=10^{-pH};"

"[H_3O^+]=10^{-3.7}=2\sdot10^{-4};"

"K_b=\dfrac{[OH^-][HCl]}{[Cl^-]};"

"[OH^-]=[HCl]=x;"

"x^2=K_b[Cl^-];"

"x^2=1\sdot10^{-4}\sdot 2\sdot10^{-4}=2\sdot10^{-4}; x=1.41\sdot10^{-2};"

"[OH^-]=1.41\sdot10^{-2};"

ii. formula for the conjugate base of "HCl" : "Cl^-;"

Answer: A. i."NH_3+H_2O\to NH_4OH;"

ii."H_2CO_3+2H_2O\to2H_3O+CO_3;"

B. i. "pH=2;"

ii. "pH=2.72;"

C. i. "[OH^-]=1.41\sdot10^{-2};"

ii. formula for the conjugate base of "HCl" is "Cl^-."